# Euler characteristic of crepant resolutions of specific modular quotient singularities

### v1

@@ 2 Preliminaries
We first consider the case when $G$ is not abelian. Then $G$ contains #$\#C(G)$ conjugacy classes of $1$ element, #$\frac{\#H-\#C(G)}{p}$ conjugacy classes of $p$ elements and $(p-1)\#C(G)$ conjugacy classes of $(H:C(G))$ elements. Hence the number of conjugacy classes of $G$ is equal to right hand side in the proposition.
On the other hand, by a corollary of Green correspondence (one can refer to [webb2016course], Corollary 11.6.5), for $G$ with its $p$-Sylow subgroup isomorphic to $C_{p}$, if one denotes number of simple $kG$-modules up to isomorphism by $l_{k}(G)$, then
#$\#\{[\text{indecomposable }kG\text{-modules}]\}=l_{k}(G)+(p-1)l_{k}(N_{G}(C_{p})).$
By Brauer’s theorem on $l_{k}(G)$, we have $l_{k}(G)=\#C(G)+\frac{\#H-\#C(G)}{p}$ and $l_{k}(N_{G}(C_{p}))=l_{k}(C(G)\times C_{p})=\#C(G)$. Therefore, the number of isomorphic classes of indecomposable $kG$-modules coincides #$\#\mathrm{Conj}(G)$.
If $G$ is abelian, then #$\#\mathrm{Conj}(G)=\#G=p\#H$, $l_{k}(G)=\#H$ and $N_{G}(C_{p})=G$. Then easy computation shows the statement again. ∎
By Corollary 2.11 and Proposition 2.14, to prove the main theorem, it suffices to show that under the assumption of the main theorem, if there exists a crepant resolution, then
@@
Note that even though we use a version of wild McKay correspondence over finite fields, we can consider the main theorem over appropriate finite fields instead, such that the $G$-actions, quotient singularties and resolutions are obtained by base change. In the remaining parts, we always assume that the order of the finite field $k=\mathbb{F}_{q}$ is large enough, such that it is a representation-theoretic splitting field of $G$ in characteristic $p$; in other words, not only Euler characteristic, but also the representation-theoretic invariants in the main theorem remain unchanged under base change to $\overline{k}$.
3 Abelian case of the main theorem
The proof starts from the case when $G$ is abelian. Now let $G\cong H\times C_{p}$ for an abelian non-modular subgroup $H$. We first give an observation on structure of such $G$.
@@ 3 Abelian case of the main theorem
Let $G\cong H\times C_{p}$ be a finite subgroup of $\mathrm{GL}(n,k)$. Then the representation space $V$ has a decomposition $V\cong\bigoplus_{i=1}^{s}V_{i}$, such that each $V_{i}$ induces an indecomposable $C_{p}$-representation of dimension $11\leqslant d_{i}\leqslant p-$ 1 and $\forall h\in H$, $h$ acts as a scalar multiplication on $V_{i}$.
@@
For $G\cong H\times C_{p}$, we can assume that $H$ consists of diagonal matrices by similarity transformation. Denote the matrix of a generator of $C_{p}$ under such transformation by $a=(a_{ij})_{i,j}$. $\forall h=\mathrm{diag}(\alpha_{1},\dots,\alpha_{n})\in H$, from $ha=ah$ we have $\alpha_{i}a_{ij}=\alpha_{j}a_{ij}$ for any $i,j$. If any $h$ in $H$ is of the form $\mathrm{diag}(c,\dots,c)$, then the lemma is trivial. So we can assume that there exists some $\alpha_{i}\neq\alpha_{j}$, and therefore $a_{ij}=a_{ji}=0$. By changing the order of basis if necessary, this implies that $a$ is a block diagonal matrix. In other words, $V$, as a representation space of $G$, can be decomposed into direct sum of subrepresentations as eigenspaces of different eigenvalues of $h$. Repeat this process on each subspace in the obtained decomposition, and then we obtain $V=\bigoplus_{t}W_{t}$, such that $\forall h\in H$ and $\forall W_{t}$, $W_{t}$ is a subspace of an eigenspace of $h$. Decompose each $W_{t}$ into direct sum of indecomposable representations of $C_{p}$. Then every components is still a $G$-representation, and the desired decomposition $V=\bigoplus_{i=1}^{s}V_{i}$ is obtained. ∎
@@
Consequently, we can choose basis of each $V_{i}$ appropriately, such that $H$ consists of diagonal matrices, and a generator of $C_{p}$ has the form $\mathrm{diag}(J_{d_{1}}(1),J_{d_{2}}(1),\dots,J_{d_{s}}(1))$, where $J_{i}(\alpha)$ denotes the Jordan block of size $i$ and eigenvalue $\alpha$.
Then we want to compute $f_{G}$ for $G\cong H\times C_{p}$ a small finite subgroup of $\mathrm{SL}(n,k)$.
For an element $L\in A(G)$, forgetting the $G$-action on it, $L$ is a Galois extension obtained by choosing a Galois extension $L_{1}/K$ with Galois group $\mathrm{Gal}(L_{1}/K)\cong H$ and an Artin-Schreier extension $L_{2}/K$, such that $L/K=L_{1}L_{2}/K$. For $L_{2}$, it depends on the choice of an element $a\in K$ such that $L_{2}=K(\alpha)$ for $\alpha$ satisfying $\wp(\alpha)=\alpha^{p}-\alpha=a$. We may choose $a$ such that the fixed generator of $C_{p}$ maps $\alpha$ to $\alpha+1$. According to [yasuda2014cyclic], Proposition 2.5, for this $G$-étale $K$-algebra, up to isomorphism, the element $a$ for the Artin-Schreier extension should be chosen from
@@
In other words, every element $L\in A(G)$ is determined by an element $L_{1}\in A(H)$ and an element $L_{2}\in A(C_{p})$ determined by $a\in K/\wp(K)\setminus\{0\}$. If we fix $L_{1}$ and the degree $-j\in\{n\in\mathbb{Z}\mid n\leqslant 0,n\notin p\mathbb{Z}_{-}\}$, then there are
@@
choices of $C_{p}$-étale extension $L_{2}$ generated by the element $\alpha$ with $v_{L_{2}}(\alpha)=-j$. $L_{2}$ is unramified if and only if $j=0$.
Fix $L\in A(G)$ and continue using notations $L_{1},L_{2},j$. We want to compute the value of $\mathbf{v}(L)$.
@@
$\mathbf{v}(L)=\mathbf{v}(L/L_{1}^{\mathrm{ur}}),$ where $\mathbf{v}(L/L_{1}^{\mathrm{ur}})$ is the value of $\mathbf{v}$ funtion defined for group $\overline{G}:=G/\mathrm{Gal}(L_{1}^{\mathrm{ur}}/K)$ over $L_{1}^{\mathrm{ur}}\cong\mathbb{F}_{q^{f}}((t))$, and $f$ is the residue degree of $L_{1}/K$.
@@
Denote the dual basis of $T=\sum\mathcal{O}_{K}x_{i}$ in $\mathrm{Hom}_{\mathcal{O}_{K}}(T,\mathcal{O}_{L})$ by $x_{1}^{*},\dots,x_{n}^{*}$. Then by [yasuda2017toward], Proposition 6.3, $\mathrm{Hom}_{\mathcal{O}_{K}}^{G}(T,\mathcal{O}_{L})$ is a free $\mathcal{O}_{K}$-module of rank $n$. If $(\sum c_{ij}x_{j}^{*})_{1\leqslant i\leqslant n}$ is an $\mathcal{O}_{K}$-basis of $\mathrm{Hom}_{\mathcal{O}_{K}}^{G}(T,\mathcal{O}_{L})$, then $\mathbf{v}(L)=\frac{1}{\#G}\mathrm{length}_{\mathcal{O}_{K}}\frac{\mathcal{O}_{L}}{(\mathrm{det}(c_{ij}))}=\frac{v_{L}(\mathrm{det}(c_{ij}))f_{L/K}}{e_{L/K}f_{L/K}}=\frac{v_{L}(\mathrm{det}(c_{ij}))}{e_{L/K}}$.
@@
By Lemma 3.1 and Lemma 3.3, together with convertibility and additivity of $\mathbf{v}$ functions, we only need to consider the case when $L_{1}$ is a cyclic totally ramified extension and the representation is indecomposable.
@@
Assume that $1<n\leqslant p$, and that $G$ is generated by $\sigma,\tau$ of the form
@@
and that $L_{1}$ is totally ramified with ramification index $l$. (We allow $G\subseteq\mathrm{GL}(n,k)$ here.)
@@
The case when $l=1$ has been computed in [yasuda2017toward], Example 6.8.
Now we assume $l>1$. Without loss of generality, we can choose the generator $\alpha$ of $L_{2}$ and a uniformizer $\beta$ of $\mathcal{O}_{L_{1}}$ such that
@@
We define formal binomial coefficients
(${\alpha\choose i}:=\frac{\alpha(\alpha-1)\dots(\alpha-i+1)}{i!}$
for $i=1,\dots,n-1$. Since $i<n\leqslant p$, they are well-defined. Additionally denote (${\alpha\choose 0}=1$. Then formal binomial coefficients satisfy
(${\alpha+1\choose i}={\alpha\choose i}+{\alpha\choose i-1}.$
Let $x_{1}^{*},\dots,x_{n}^{*}$ be the dual basis of $T=\sum\mathcal{O}_{K}x_{i}$ in $\mathrm{Hom}_{\mathcal{O}_{K}}(T,\mathcal{O}_{L})\hookrightarrow\mathrm{Hom}_{K}(\sum Kx_{i},L)$. Then the action of $\sigma$ and $\tau$ on $T$ is given by
@@
$\begin{pmatrix}\beta x_{1}^{*}&\cdots&\beta x_{n}^{*}\end{pmatrix}\begin{pmatrix}0&0&\cdots&0&1\\ 0&0&\cdots&1&{\alpha\choose 1}\\ \vdots&\vdots&\iddots&\vdots&\vdots\\ 0&1&\cdots&\cdots&{\alpha\choose n-2}\\ 1&{\alpha\choose 1}&\cdots&{\alpha\choose n-2}&{\alpha\choose n-1}\\ \end{pmatrix}$
over $K$. Denote this basis by $(\phi_{1},\dots,\phi_{n})=\beta(x_{1}^{*},\dots,x_{n}^{*})\mathbf{A}$, and the matrix $\mathbf{A}$ has determinant ±$\pm 1$. If we take
@@
Now we goes back to general case of the abelian group $G\cong H\times C_{p}$ and the $G$-étale $K$-algebra $L\in A(G)$.
@@
$\mathrm{age}(L_{1}):=\sum_{i=1}^{s}d_{i}(\frac{1}{l_{i}}-\left\lfloor\frac{1}{l_{i}}\right\rfloor).$
@@
When $H=\{e\}$, $\mathrm{Sht}^{L}_{V}(j)=\mathrm{sht}_{V}(j)+n-s$, where $\mathrm{sht}_{V}$ is the shift number defined in [yasuda2014cyclic].
@@
Note that $l_{i}$ and $\mathrm{age}(L_{1})$ are totally determined by $L_{1}$. For the part of generalized upper shift numbers, it is determined by $L_{1}$ and $j$. We also have the following property by easy computation.
@@
Let $G\cong H\times C_{p}$. If there exists a crepant resolution $Y\to X$, then $D_{V}=p$, and every $f_{G^{\prime}}$ with $G^{\prime}\cong H^{\prime}\times C_{p}\subseteq G$ converges as a polynomial, with $S(f_{G^{\prime}})=(p^{2}-1)S(f_{H^{\prime}}).$
@@
Therefore, $f_{G}(q)$ converges if and only if $D_{V}\geqslant p$; if so, we can write
@@
$=$ $\sum_{L_{1}\in A(H)}q^{n-\mathrm{age}(L_{1})}(p-1+p\frac{1-q^{-1}}{1-q^{p-1-D_{V}}}\sum_{r=1}^{p-1}q^{r-\mathrm{Sht}^{L}_{V}(r)}).$
@@
Since $D_{V}$ is equal for all $G^{\prime}\cong H^{\prime}\times C_{p}\subseteq G$, we complete the proof of this proposition.
@@
As a byproduct of this proposition, if $D_{V}\geqslant p$, then the quotient singularity $X$ is log terminal because of the convergence of stringy motif. The reader can refer to [yasuda2014cyclic] for details in $p$-cyclic McKay correspondence.
@@
Main theorem holds when $G$ is abelian.
@@
By Proposition 3.9, if there exists a crepant resolution, then
@@
4 Non-abelian case of the main theorem
To study non-abelian case, we need the following lemma.

### v2

@@ 2 Preliminaries
We first consider the case when $G$ is not abelian. Then $G$ contains #C(G)\#C(G) conjugacy classes of one element, (#H#C(G))/p(\#H-\#C(G))/{p} conjugacy classes of $p$ elements and (p1)#C(G)(p-1)\#C(G) conjugacy classes of $[H:C(G)]$ elements. Hence the number of conjugacy classes of $G$ is equal to the right-hand side of the proposition.
On the other hand, by a corollary of Green correspondence (one can refer to [12], Corollary 11.6.5), for $G$ with its $p$-Sylow subgroup isomorphic to $C_{p}$, if one denotes the number of simple $kG$-modules up to isomorphism by $l_{k}(G)$, then
$\mathrm{Ind}_{k}(G)=l_{k}(G)+(p-1)l_{k}(N_{G}(C_{p})).$
By Brauer’s theorem on $l_{k}(G)$, we have $l_{k}(G)=\#C(G)+(\#H-\#C(G))/{p}$ and $l_{k}(N_{G}(C_{p}))=l_{k}(C(G)\times C_{p})=\#C(G)$. Therefore, the number of isomorphism classes of indecomposable $kG$-modules coincides with $\mathrm{Conj}(G)$.
If $G$ is abelian, then $\mathrm{Conj}(G)=\#G=p\#H$, $l_{k}(G)=\#H$ and $N_{G}(C_{p})=G$. Then an easy computation shows the statement again. ∎
In the last part of this section, we introduce the main idea for the proof of our main result. By Corollary 2.16 and Proposition 2.19, to prove the main theorem, it suffices to show that under the assumption of the main theorem, if there exists a crepant resolution, then
@@
Note that even though we use a version of the wild McKay correspondence over finite fields, we can consider the main theorem over some appropriate finite fields instead, such that the $G$-actions, quotient singularities and resolutions are obtained by base change. In the remaining parts, we always assume that the order of the finite field $k=\mathbb{F}_{q}$ is large enough, such that it is a representation-theoretic splitting field of $G$ in characteristic $p$; in other words, not only the Euler characteristic, but also the representation-theoretic invariants in the main theorem remain unchanged under the base change to $\overline{k}$.
3 The abelian case of the main theorem
The proof starts from the case when $G$ is abelian. Now let $G\cong H\times C_{p}$ for an abelian non-modular subgroup $H$. We first give an observation on the structure of such $G$.
@@ 3 The abelian case of the main theorem
Let $G\cong H\times C_{p}$ be a finite subgroup of $\mathrm{GL}(n,k)$. Then the representation space $V$ has a direct sum decomposition $V\cong\bigoplus_{i=1}^{s}V_{i}$, such that each $V_{i}$ induces an indecomposable $C_{p}$-representation of dimension $1\leqslant d_{i}\leqslant p-1$ and $\forall h\in H$, $h$ acts on $V_{i}$ as a scalar multiplication.
@@
For $G\cong H\times C_{p}$, we can assume that $H$ consists of diagonal matrices by similarity transformation. Denote the matrix of a generator of $C_{p}$ under such transformation by $a=(a_{ij})_{i,j}$. $\forall h=\mathrm{diag}(\alpha_{1},\dots,\alpha_{n})\in H$, from $ha=ah$ we have $\alpha_{i}a_{ij}=\alpha_{j}a_{ij}$ for any $i,j$. If any $h$ in $H$ is of the form $\mathrm{diag}(c,\dots,c)$, then the lemma is trivial. So we can assume that there exists some $\alpha_{i}\neq\alpha_{j}$, and therefore $a_{ij}=a_{ji}=0$. By changing the order of the basis if necessary, this implies that $a$ is a block diagonal matrix. In other words, $V$, as a representation space of $G$, can be decomposed into a direct sum of subrepresentations as eigenspaces of different eigenvalues of $h$. Repeat this process on each subspace in the obtained decomposition for all $h\in H$, and then we obtain $V=\bigoplus_{t}W_{t}$, such that $\forall h\in H$ and $\forall W_{t}$, $W_{t}$ is a subspace of an eigenspace of $h$. Decompose each $W_{t}$ into the direct sum of indecomposable representations of $C_{p}$. Then every summand is still a $G$-representation, and the desired decomposition $V=\bigoplus_{i=1}^{s}V_{i}$ is obtained. ∎
@@
Consequently, we can choose a basis of each $V_{i}$ appropriately, such that $H$ consists of diagonal matrices, and a generator of $C_{p}$ has the form $\mathrm{diag}(J_{d_{1}}(1),J_{d_{2}}(1),\dots,J_{d_{s}}(1))$, where $J_{i}(\alpha)$ denotes the Jordan block of size $i$ and eigenvalue $\alpha$.
Then we want to compute $f_{G}$ for $G\cong H\times C_{p}$, a small finite subgroup of $\mathrm{SL}(n,k)$.
For an element $L\in A(G)$, forgetting the $G$-action on it, $L$ is a Galois extension obtained by choosing a Galois extension $L_{1}/K$ with Galois group $\mathrm{Gal}(L_{1}/K)\cong H$ and an Artin-Schreier extension $L_{2}/K$, such that $L/K=L_{1}L_{2}/K$. For $L_{2}$, it depends on the choice of an element $a\in K$ such that $L_{2}=K(\alpha)$ for some $\alpha$ satisfying $\wp(\alpha)=\alpha^{p}-\alpha=a$. We may choose $\alpha$ such that the fixed generator of $C_{p}$ maps $\alpha$ to $\alpha+1$. According to [17], Proposition 2.5, for this $C_{p}$-étale $K$-algebra, up to isomorphism, the element $a$ for the Artin-Schreier extension can be chosen from
@@
In other words, every element $L\in A(G)$ is determined by an element $L_{1}\in A(H)$ and an element $L_{2}\in A(C_{p})$ with $a\in K/\wp(K)\setminus\{0\}$. If we fix $L_{1}$ and the degree $j\in\{n\in\mathbb{Z}\mid n\leqslant 0,n\notin p\mathbb{Z}_{-}\}$, then there are
@@
choices of the $C_{p}$-étale extension $L_{2}$ generated by the element $\alpha$ with $v_{L_{2}}(\alpha)=-j$. $L_{2}$ is unramified if and only if $j=0$.
Fix $L\in A(G)$ and continue using the notations $L_{1},L_{2},j$. We want to compute the value of $\mathbf{v}(L)$.
@@
$\mathbf{v}(L)=\mathbf{v}(L/L_{1}^{\mathrm{ur}})$. Here $L_{1}^{\mathrm{ur}}$ is the maximal unramified extension of $L_{1}$, $\mathbf{v}(L/L_{1}^{\mathrm{ur}})$ is the value of the $\mathbf{v}$ function defined for the group $\overline{G}:=G/\mathrm{Gal}(L_{1}^{\mathrm{ur}}/K)$ over $L_{1}^{\mathrm{ur}}\cong\mathbb{F}_{q^{f}}((t))$, and $f$ is the residue degree of $L_{1}/K$.
@@
Denote the dual basis of $T=\sum\mathcal{O}_{K}x_{i}$ in $\mathrm{Hom}_{\mathcal{O}_{K}}(T,\mathcal{O}_{L})$ by $x_{1}^{*},\dots,x_{n}^{*}$. Then by [18], Proposition 6.3, $\mathrm{Hom}_{\mathcal{O}_{K}}^{G}(T,\mathcal{O}_{L})$ is a free $\mathcal{O}_{K}$-module of rank $n$. If $(\sum c_{ij}x_{j}^{*})_{1\leqslant i\leqslant n}$ is an $\mathcal{O}_{K}$-basis of $\mathrm{Hom}_{\mathcal{O}_{K}}^{G}(T,\mathcal{O}_{L})$, then
$\mathbf{v}(L)=\frac{1}{\#G}\mathrm{length}_{\mathcal{O}_{K}}\frac{\mathcal{O}_{L}}{(\mathrm{det}(c_{ij}))}=\frac{v_{L}(\mathrm{det}(c_{ij}))f_{L/K}}{e_{L/K}f_{L/K}}=\frac{v_{L}(\mathrm{det}(c_{ij}))}{e_{L/K}}.$
@@
By Lemma 3.20 and Lemma 3.22, together with the convertibility and additivity of $\mathbf{v}$ functions, we only need to consider the case when $L_{1}$ is a cyclic totally ramified extension and the representation is indecomposable.
@@
Assume that $1\leqslant n\leqslant p$, and that $G$ is generated by $\sigma,\tau$ of the form
@@
and that $L_{1}/K$ is totally ramified with ramification index $l$. (We allow $G\subseteq\mathrm{GL}(n,k)$ here.)
@@
For the case when $l=1$, the value of the $\mathbf{v}$ function has been computed in [18], Example 6.8.
Now we assume that $n,l>1$. Without loss of generality, we can choose the generator $\alpha$ of $L_{2}$ and a uniformizer $\beta$ of $\mathcal{O}_{L_{1}}$ such that
@@
We define the formal binomial coefficients
$\binom{\alpha}{i}:=\frac{\alpha(\alpha-1)\dots(\alpha-i+1)}{i!}$
for $i=1,\dots,n-1$. Since $i<n\leqslant p$, they are well-defined. Additionally denote $\binom{\alpha}{0}=1$. Then the formal binomial coefficients satisfy
$\binom{\alpha+1}{i}=\binom{\alpha}{i}+\binom{\alpha}{i-1}$.
Let $x_{1}^{*},\dots,x_{n}^{*}$ be the dual basis of $T=\sum\mathcal{O}_{K}x_{i}$ in $\mathrm{Hom}_{\mathcal{O}_{K}}(T,\mathcal{O}_{L})\hookrightarrow\mathrm{Hom}_{K}(\sum Kx_{i},L)$. Then the actions of $\sigma$ and $\tau$ on $T$ are given by
@@
$\begin{pmatrix}\beta x_{1}^{*}&\cdots&\beta x_{n}^{*}\end{pmatrix}\begin{pmatrix}0&0&\cdots&0&1\\ 0&0&\cdots&1&\binom{\alpha}{1}\\ \vdots&\vdots&\iddots&\vdots&\vdots\\ 0&1&\cdots&\cdots&\binom{\alpha}{n-2}\\ 1&\binom{\alpha}{1}&\cdots&\binom{\alpha}{n-2}&\binom{\alpha}{n-1}\\ \end{pmatrix}$
over $K$. Denote this basis by $(\phi_{1},\dots,\phi_{n})=\beta(x_{1}^{*},\dots,x_{n}^{*})\mathbf{A}$, and then the matrix $\mathbf{A}$ has determinant $\pm 1$. If we take
@@
For the case when $n=1$, it is easy to show the lemma with a similar statement only considering the tame part.
Now we go back to the general case for the abelian group $G\cong H\times C_{p}$ and the GG-étale $K$-algebra $L\in A(G)$.
@@
$\mathrm{age}(L_{1}):=\sum_{i=1}^{s}d_{i}\left(\frac{1}{l_{i}}-\left\lfloor\frac{1}{l_{i}}\right\rfloor\right).$
@@
If $H=\{e\}$, then $\mathrm{Sht}^{L}_{V}(j)=\mathrm{sht}_{V}(j)+n-s$, where $\mathrm{sht}_{V}$ is the shift number defined in [17]. If $d_{i}=1$ for any $i$, then $\mathrm{age}(L_{1})=\mathrm{age}(g)$, where $g\in G$ satisfies that for any $N_{i}=L_{1}^{\mathrm{nr}}(t_{i})$, $g(t_{i})=\zeta_{l_{i}}t_{i},$ and $\mathrm{age}(g)$ is the age grading defined in [9].
@@
Note that $l_{i}$ and $\mathrm{age}(L_{1})$ are completely determined by $L_{1}$. For the part of generalized upper shift numbers, it is determined by $L_{1}$ and $j$. We also have the following property by an easy computation.
@@
Let $G\cong H\times C_{p}$. If there exists a crepant resolution $Y\to X$, then $D_{V}=p$, and every $f_{G^{\prime}}$ with $G^{\prime}\cong H^{\prime}\times C_{p}\subseteq G$ is a polynomial in $q$, with $S(f_{G^{\prime}})=(p^{2}-1)S(f_{H^{\prime}}).$
@@
Therefore, $f_{G}(q)$ is a rational function if and only if $D_{V}\geqslant p$; if so, we can write
@@
$=$ $\sum_{L_{1}\in A(H)}q^{n-\mathrm{age}(L_{1})}\left(p-1+p\frac{1-q^{-1}}{1-q^{p-1-D_{V}}}\sum_{r=1}^{p-1}q^{r-\mathrm{Sht}^{L}_{V}(r)}\right).$
@@
Since $D_{V}$ is equal for all $G^{\prime}\cong H^{\prime}\times C_{p}\subseteq G$, we complete the proof of this proposition by similar computations.
@@
As a byproduct of this proposition, if $D_{V}\geqslant p$, then the quotient singularity $X$ is log terminal because of the convergence of its stringy motif. The reader can refer to [17] for details in the $p$-cyclic McKay correspondence.
@@
Our main theorem holds when $G$ is abelian.
@@
By Proposition 2.17 and Proposition 3.28, if there exists a crepant resolution, then
@@
4 The non-abelian case of the main theorem
To study the non-abelian case, we need the following lemma.